OpenStudy (anonymous):

When this polynomial is simplified, what is the coefficient of x^2: 10x^2+8x-6x^3+4x^2+12x^3-x+20

jimthompson5910 (jim_thompson5910):

10x^2+4x^2 = (10+4)x^2 = 14x^2 So the answer is 14